If the source charge is +4.31x10^3 C and the test charge is placed 7.11x10^-12m away from it?


1
Expert's answer
2021-03-23T11:13:03-0400

... then the electric field is


"E=k\frac{q}{r^2}=9\cdot10^9\cdot\frac{4.31\cdot10^{3}}{(7.11\cdot10^{-12})^2}=7.67\cdot10^{35}\ (V/m)" . Answer